# Algorithmic Collusion under Observed Demand Shocks

### v3

@@
(November 3, 2025)
This paper investigates how the observability of demand shocks influences pricing patterns and market outcomes when firms delegate pricing decisions to Q-learning algorithms. Simulations show that demand observability induces Q-learning agents to adapt prices to demand fluctuations, giving rise to distinctive demand-contingent pricing patterns across the discount factor $\delta$, consistent with Rotemberg and Saloner (1986). When $\delta$ is high, they learn procyclical pricing, charging higher prices in higher demand states. In contrast, at low $\delta$, they lower prices during booms and raise them during downturns, exhibiting countercyclical pricing. Moreover, Q-learning agents autonomously learn to sustain supracompetitive profits. I further examine how the information available to algorithms shapes their learned pricing patterns and find that price memory is indispensable for generating countercyclical pricing. These findings suggest that Q-learning algorithms, through pure trial and error and without prior knowledge of the environment, capture the stronger deviation incentives during booms and the discount factor’s role in balancing short-term gains against long-term continuation values, thereby reproducing the cyclicality of pricing patterns predicted by collusion theory.
@@ 1 Introduction
Second, Q-learning agents can autonomously learn to sustain supracompetitive profits under observed demand shocks. Relative to the fixed-demand benchmark, learning under observed demand shocks exhibits a profit reversal—yielding higher profits when $\delta$ is low but lower profits when $\delta$ is high. [^8] The fixed-demand benchmark is constructed by training Q-learning agents separately in each fixed demand state and averaging profits weighted by the demand distribution. This profit reversal may arise because, under observed demand shocks, algorithms are exposed to payoff variation across demand states, generating heterogeneous continuation values. When $\delta$ is low, fixed-demand learning drives prices in both demand states toward the competitive level. Under observed demand shocks, however, continuation values occasionally carried over from booms slow the downward adjustment of prices in downturns, thereby sustaining higher prices and raising expected profits. As $\delta$ increases, greater weight on continuation values makes high prices more sustainable. In the fixed-demand environment, prices eventually converge to a single high price. Under observed demand shocks, however, coordinating demand-contingent pricing introduces frictions between agents, leading to asymmetric pricing within the price cycle that, although infrequent, erodes profits.
@@
Overall, this paper bridges classic collusion theory and modern reinforcement learning by showing that Q-learning algorithms, when exposed to observed demand shocks, experience payoff variation across demand states and prices, thereby autonomously reproducing the cyclicality of pricing patterns predicted by Rotemberg and Saloner (1986). These learning outcomes suggest that, despite having no explicit understanding of equilibrium or intertemporal incentives, the algorithms, through pure trial and error, come to capture the key insights underlying this cyclicality, namely that (i) stronger deviation incentives arise during booms and (ii) the trade-off between short-term gains and long-term continuation values is governed by $\delta$, and behave as if guided by them.
@@ 3.3 Hyperparameters
I also simulate environments in which the demand state is fixed at either $L$ or $H$. Learning in these fixed-demand settings provides a natural benchmark for comparison with learning under observed demand shocks. This benchmark reflects practical situations in which firms train pricing algorithms in static environments before deploying them in dynamic markets subject to stochastic demand shocks.
@@
Although convergence of Q-learning has been proven in single-agent settings under specific conditions (Watkins and Dayan, 1992), it is not theoretically guaranteed in strategically interdependent environments. Following Calvano et al. (2020), I employ an empirical convergence criterion: learning is considered complete when each agent’s optimal strategy remains unchanged for $100,000$ consecutive periods. [^19] Specifically, convergence is achieved when, for each agent $i$ and state $s$, the optimal price $p_{it}(s)=\underset{p\in A}{\mathrm{argmax}}\,\,Q_{i}(s,p)$ remains constant over $100,000$ consecutive iterations. The simulation terminates once this criterion is satisfied or after one billion iterations, whichever occurs first. Under the baseline hyperparameters $\alpha=0.15$ and $\beta=4\times 10^{-6}$, the average number of iterations required for convergence is 2,677,4362,677,436.202020Convergence under fixed demand states $L$ and $H$ requires $1,642,801$ and $1,815,529$ iterations, respectively. These smaller numbers are expected, as the state space in each single-demand simulation is only half the size of the state space under observed demand shocks.
@@ 4.4 Profit Dynamics
A possible explanation for this profit reversal is as follows. When $\delta$ is low, Q-learning agents in the fixed-demand environment behave myopically, placing excessive weight on immediate rewards. After suffering short-term losses from being undercut, they repeatedly lower prices, a self-reinforcing dynamic that pushes prices toward the competitive level and leaves profits persistently low. In contrast, learning under observed demand shocks exposes agents to payoff variation across demand states. Because the demand state switches stochastically, Q-updates under $L$ incorporate continuation values from $H$ about half of the time, which are typically higher than those associated with $L$. Thus, the difference in continuation values across demand states slows the downward adjustments of Q-values after being undercut in $L$, allowing agents to sustain relatively higher prices and thereby generating countercyclical pricing. This is the key difference from the fixed-demand environment, where continuation values are constructed only from payoffs generated under the same demand condition. As a result, learning under observed demand shocks yields higher profits than fixed-demand learning when $\delta$ is low.
As $\delta$ increases, agents place greater weight on continuation values and become more tolerant of temporary losses, allowing higher prices to be sustained. Under the fixed-demand environment, the absence of cross-demand disturbances facilitates convergence to a single high price. By contrast, when demand shocks are observed, payoff variation across demand states drives learning toward demand-contingent pricing. However, this process inevitably introduces coordination frictions and generates asymmetric pricing. With greater tolerance, such asymmetries persist within the price cycle, eroding profits and ultimately making learning under observed demand less profitable than fixed-demand learning when $\delta$ is high.
@@ 5.3 Summary
First, since the state variable contains rich information, agents experience payoff variation across demand states and prices. Through repeated interactions, they gradually recognize that profits are higher under $H$, and that being undercut in this demand state results in a larger immediate payoff loss, leading to larger downward adjustments in Q-values. This learning process effectively captures the stronger deviation incentives during booms.
Second, when $\delta$ is high, the greater weight on continuation values allows high prices to be sustained across both demand states. Through exploration, agents learn that charging high prices under $H$ yields greater immediate rewards and correspondingly higher Q-values than rigid pricing, leading them to set higher prices under $H$ relative to $L$. Although deviation incentives are stronger under $H$, the greater weight on continuation values makes prices more resistant to short-term deviations and thereby generates procyclical pricing. As $\delta$ decreases, greater emphasis is placed on short-term gains. Given that high prices under $H$ are more vulnerable to undercutting than those under $L$, agents learn to lower prices under $H$ while maintaining relatively high prices under $L$, thereby generating countercyclical pricing.
In this sense, even though Q-learning agents have no explicit knowledge of these two core insights, through pure trial and error they come to behave as if guided by them, thereby reproducing the pricing patterns predicted by collusion theory.
@@ 7 Conclusion
This paper examines how the observability of demand shocks shapes the learning outcomes of pricing algorithms. The simulation results show that Q-learning agents autonomously adapt to demand fluctuations and develop demand-contingent pricing, which is procyclical at high $\delta$ and countercyclical at low $\delta$, consistent with Rotemberg and Saloner (1986). They also sustain supracompetitive profits, demonstrating the robustness of algorithmic collusion under observed demand shocks. Further analysis reveals that the learned pricing patterns critically depend on the information embedded in the state variable. Removing demand memory weakens demand-contingent pricing and drives learning toward rigid pricing. Price memory has a more complex effect: at high $\delta$, agents learn procyclical pricing regardless of whether price memory is present, whereas at low $\delta$, the absence of price memory leads to uniformly rigid pricing, making it essential for sustaining countercyclical pricing. These findings shed light on how Q-learning agents generate distinctive demand-contingent pricing patterns across $\delta$. Through repeated exposure to payoff variation across demand states and prices, they capture both the stronger deviation incentives that arise during booms and the discount factor’s role in balancing short-term gains and long-term continuation values, thereby reproducing the pricing patterns predicted by collusion theory.
@@
This paper also uncovers an intriguing yet puzzling profit reversal: learning under observed demand shocks yields higher profits than the fixed-demand benchmark when $\delta$ is low but lower profits when $\delta$ is high. One possible explanation is that exposure to payoff variation across demand states alters how agents, at different discount factors, evaluate the trade-off between short-term gains and long-term continuation values. Further simulation evidence is needed to assess its robustness, and a formal theoretical framework explaining this reversal would substantially deepen our understanding of algorithmic learning dynamics.

### v4

@@
(December 7, 2025)
This paper examines how the observability of demand shocks influences pricing patterns and market outcomes when firms delegate pricing decisions to Q-learning algorithms. Simulations show that demand observability induces Q-learning agents to adapt prices to demand fluctuations, giving rise to distinctive demand-contingent pricing patterns across the discount factor $\delta$, consistent with Rotemberg and Saloner (1986). When $\delta$ is high, they learn procyclical pricing, charging higher prices in higher demand states. In contrast, at low $\delta$, they lower prices during booms and raise them during downturns, exhibiting countercyclical pricing. Q-learning agents also autonomously sustain supracompetitive profits, indicating that demand observability does not hinder algorithmic collusion. I further explore how the information available to algorithms shapes their learned pricing behavior. Overall, the results suggest that, through pure trial and error, Q-learning algorithms internalize both the stronger deviation incentives during booms and the trade-off between short-term gains and long-term continuation values governed by the discount factor, thereby reproducing the cyclicality of pricing patterns predicted by collusion theory.
@@ 1 Introduction
Second, Q-learning agents can autonomously learn to sustain supracompetitive profits under observed demand shocks, indicating that demand observability does not hinder algorithmic collusion. However, relative to the fixed-demand benchmark, learning under observed demand shocks exhibits a profit reversal: it delivers higher profits when $\delta$ is low but lower profits when $\delta$ is high. [^8] The fixed-demand benchmark is constructed by training Q-learning agents separately in each fixed demand state and averaging profits weighted by the demand distribution. This profit reversal arises from the demand-contingent pricing, highlighting that the observability of demand shocks is a double-edged sword for profitability.
@@
Overall, this paper bridges classic collusion theory and modern reinforcement learning by showing that Q-learning algorithms, when exposed to observed demand shocks, experience payoff variation across demand states and prices, thereby autonomously reproducing the cyclicality of pricing patterns predicted by Rotemberg and Saloner (1986). These learning outcomes suggest that, despite having no explicit understanding of equilibrium or intertemporal incentives, the algorithms, through pure trial and error, come to internalize the key insights underlying this cyclicality, namely that (i) stronger deviation incentives arise during booms and (ii) the trade-off between short-term gains and long-term continuation values is governed by $\delta$, and behave as if guided by them.
@@ 3.3 Hyperparameters
I also simulate environments in which the demand state is fixed at either $L$ or HH. [^19] In the fixed-demand simulations, the demand state is constant, so the state variable reduces to $s_{t}=(p_{1t-1},p_{2t-1})$. Learning in these fixed-demand settings provides a natural benchmark for comparison with learning under observed demand shocks. This benchmark reflects practical situations in which firms train pricing algorithms in static environments before deploying them in dynamic markets subject to stochastic demand shocks.
@@
Although convergence of Q-learning has been proven in single-agent settings under specific conditions (Watkins and Dayan, 1992), it is not theoretically guaranteed in strategically interdependent environments. Following Calvano et al. (2020), I employ an empirical convergence criterion: learning is considered complete when each agent’s optimal strategy remains unchanged for $100,000$ consecutive periods. [^20] Specifically, convergence is achieved when, for each agent $i$ and state $s$, the optimal price $p_{it}(s)=\underset{p\in A}{\mathrm{argmax}}\,\,Q_{i}(s,p)$ remains constant over $100,000$ consecutive iterations. The simulation terminates once this criterion is satisfied or after one billion iterations, whichever occurs first. Under the baseline hyperparameters $\alpha=0.15$ and $\beta=4\times 10^{-6}$, the average number of iterations required for convergence is 2,677,4362,677,436.212121Convergence under fixed demand states $L$ and $H$ requires $1,642,801$ and $1,815,529$ iterations, respectively. These smaller numbers are expected, as the state space in each single-demand simulation is only half the size of the state space under observed demand shocks.
@@ 4.4 Profit Dynamics
Thus, relative to fixed-demand learning, the profit reversal arises naturally from the demand-contingent pricing learned under observed demand shocks, highlighting that demand observability is a double-edged sword for profitability.
@@ 5.3 Summary
When $\delta$ is low, immediate gains dominate the learning process. Through exploration, Q-learning agents discover that undercutting the rival under $H$ yields higher Q-values for lower-price actions than under $L$, as immediate deviation gains are typically much larger under HH. [^32] Q-updates also include realized discounted continuation values. However, these continuation-value adjustments are similar on average for deviations originating from both demand states. As demand shocks are i.i.d. and Q-learning operates with bounded memory, deviations from $H$ and $L$ lead to indistinguishable post-deviation price dynamics. As a result, agents cannot learn harsher punishments for deviations under $H$. Since continuation values matter little when $\delta$ is low, charging the same price cannot sustain Q-values as high as those obtained from undercutting. As a result, agents are inclined to undercut each other under $H$, driving prices toward the competitive level. By contrast, undercutting under $L$ is less attractive, allowing agents to maintain relatively higher prices under $L$. This gives rise to countercyclical pricing.
When $\delta$ is high, continuation values play a dominant role in Q-updates, making high prices easier to sustain in both demand states. Through exploration, agents learn to charge higher prices under $H$, because doing so generates higher Q-values. Although deviation incentives remain stronger under $H$, the greater weight placed on continuation values offsets the immediate gains from undercutting, making coordinated high prices more resistant to short-run deviations. This generates procyclical pricing.
In this sense, even though Q-learning agents have no explicit knowledge of these two core insights, their Q-values encode them through pure trial and error. As a result, agents come to behave as if guided by these insights, thereby reproducing the pricing patterns predicted by collusion theory.
@@ 7 Conclusion
This paper examines how the observability of demand shocks shapes the learning outcomes of pricing algorithms. The simulation results show that Q-learning agents autonomously adapt to demand fluctuations and develop demand-contingent pricing, which is procyclical at high $\delta$ and countercyclical at low $\delta$, consistent with Rotemberg and Saloner (1986). They also sustain supracompetitive profits, demonstrating the robustness of algorithmic collusion under observed demand shocks. Further analysis reveals that the learned pricing patterns critically depend on the information embedded in the state variable. Removing demand memory weakens demand-contingent pricing and drives learning toward rigid pricing. Price memory has a more complex effect: at high $\delta$, agents learn procyclical pricing regardless of whether price memory is present, whereas at low $\delta$, the absence of price memory leads to uniformly rigid pricing, making it essential for sustaining countercyclical pricing. These findings shed light on how Q-learning agents generate distinctive demand-contingent pricing patterns across $\delta$. Through repeated exposure to payoff variation across demand states and prices, they come to internalize both the stronger deviation incentives that arise during booms and the discount factor’s role in balancing short-term gains and long-term continuation values, thereby reproducing the pricing patterns predicted by collusion theory.
@@
This paper also uncovers a profit reversal: learning under observed demand shocks yields higher profits than the fixed-demand benchmark when $\delta$ is low but lower profits when $\delta$ is high. This highlights that demand observability is a double-edged sword for profitability. A promising direction for future research is developing a formal framework that characterizes the learning dynamics underlying this reversal and exploring how it extends to richer information environments or more complex market structures.